# The Kalb-Ramond field and Gravitational Parity Violation

### v1

@@ Abstract
Gravitational wave physics can probe theories that extend beyond General Relativity. Motivated by recent attention on the Kalb-Ramond field as a dark matter candidate, in this work, we study a parity violating dimension four operator which couples the dual Riemann curvature to the 2-form field. After mapping the equations of motion for the right- and left-handed gravitational wave amplitudes to the novel parameterization recently presented in Jenks:2023pmk , we discuss various constraints on the model parameters in light of the coincident electromagnetic/gravitational wave signal of GW170817 and the GWTC-3 dataset.
@@ I Introduction
Dynamical Chern-Simons theory (dCS) is likely the most studied parity violating model for gravity. In addition to the two usual massless spin-2 degrees of freedom in GR, dCS includes a pseudoscalar field $\phi$ with a canonical kinetic term coupled to the Pontryagin density $\tilde{R}_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, where $\tilde{R}_{\mu\nu\rho\sigma}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}R^{\alpha\beta}_{\ \ \ \rho\sigma}$ is the dual Riemann tensor, which arises in many contexts Yunes:2009hc (see Alexander:2009tp and references therein for a review). Interesting physics follows if the dCS pseudoscalar acts as the inflaton. This includes a preheating mechanism Alexander:2014bsa , leptogenesis Alexander:2004us , as well as the potential for parity violation to be detected in large-scale structure via galaxy correlations Creque-Sarbinowski:2023wmb . Perturbations to rotating black holes were also studied in dCS theory recently in Wagle:2023fwl , and in Nojiri:2019nar , the propagation of gravitational waves in dCS were compared to standard $f(R)$ modified gravity theories.
There are numerous other parity violating theories on the market. One such class is the ghost-free higher derivative models Crisostomi:2017ugk ; Qiao:2021fwi , which are characterized by Lagrangians of the form $\mathcal{L}(g,\partial g,\partial^{2}g)$, and individual terms are contractions of the metric, Riemann, and the Levi-Civita tensors. Higher derivative theories are generically plagued by Ostrogradsky instabilities Woodard:2015zca , therefore such theories can at best be valid up to the mass scale of the ghost degrees of freedom from an effective field theory standpoint. Adjacent to these are the ghost-free scalar tensor models, such as those discussed in Nishizawa:2018srh . Here, derivatives of a scalar field couple to contractions of the Riemann and Levi-Civita tensors. As in the higher order derivative theories, constraints on the Lagrangian coefficients must be imposed in order to avoid Ostrogradsky ghosts. There is also a parity violating addition to the Hor̂ava-Lifshitz approach to quantum gravity Horava:2009uw ; Wang:2017brl , which was studied in the context of gravitational wave physics in Zhu:2013fja . A different construction was proposed in Takahashi:2022mew ; here, the authors employed conformal transformations of the metric where the conformal factor is a function of the Pontryagin density. This produces a parity violating effect in the Einstein frame, which was recently used to study modifications of the Kerr metric Tahara:2023pyg .
@@
where the right- and left-handed strains are related to the usual +,×+$\times$+ , × amplitudes by $h_{\text{R,L}}=\frac{1}{\sqrt{2}}(h_{+}\pm ih_{\times}).$ Varying the Einstein-Hilbert Lagrangian with respect to each amplitude results in simple wave equations for each. Parity violation will then follow provided an additional interaction term sources the left- and right-handed amplitudes with a different sign, say
@@
Parity violating effects in the gravitational sector generically demand a non-minimal coupling to matter, such as the pseudoscalar in dCS theory. It is interesting to entertain the possibility that such matter belongs to a dark sector. Very recently, the Kalb-Ramond (KR) two-form field $B_{\mu\nu}$ was studied as a dark matter candidate Capanelli:2023uwv . The authors showed that the KR field and Kalb-Ramond-like-particles (KRLPs) can be produced via freeze-in and freeze-out mechanisms and can account for the relic density of dark matter today. Others have considered the KR field as a portal between the SM and dark sectors Dashko:2018dsw ; Dick:2020eop , and more generally in early universe scenarios BeltranAlmeida:2018nin ; Almeida:2019xzt including leptogenesis deCesare:2014dga , as well as late time cosmic acceleration Orjuela-Quintana:2022jrg ; BeltranAlmeida:2019fou .
@@ II Features of the KR Lagrangian
Thus although $A_{\mu}$ has four components, only three are independent. The subsidiary gauge transformation (10) is a shift involving an exact 1-form.
The gauge invariance (9), (10) is broken explicitly by the potential $V$ and the interaction (6). This is analogous to the Proca theory of a massive vector boson, which is known to be dual to the massive KR field in certain regimes Hell:2021wzm ; Capanelli:2023uwv . Regarding interactions with the gravitational waves, we will see that a simple and interesting scenario occurs when the KR field’s magnitude settles at a non-zero, constant (classical) value, $B^{2}=B_{(0)}^{2}$. This can be modeled by a potential of the form
@@
where $\bar{g}^{\mu\nu}$ is the unperturbed metric. The explicit form of the potential (11) is not relevant for this work, beyond requiring that it possesses a nontrivial stable minimum. In theories of this type, such potentials have the consequence that the physics can display (spontaneous) Lorentz violation, despite the Lagrange density being Lorentz invariant Altschul:2009ae .
@@ V GW solution in FLRW background
where primes denote partial derivatives with respect to conformal time $\eta$, $\mathcal{H}\equiv a^{\prime}/a$ is the Hubble parameter, $\lambda_{\text{R,L}}=\pm 1,$ and $\Lambda_{\text{PV}}$ is the EFT cutoff. The summation indices are such that $n$ runs over odd integers and $m$ runs over even integers, and the $\{\alpha_{n},\beta_{n},\gamma_{m},\delta_{m}\}$ functions generically depend on the conformal time and are theory dependent. For metric fluctuations on top of an FLRW cosmology in conformal coordinates, the equations of motion for the gravitational wave amplitudes in our model become
$\begin{split}h_{\text{L}}^{\prime\prime}+2\mathcal{H}h_{\text{L}}^{\prime}-\partial^{2}_{z}h_{\text{L}}-\frac{\xi}{m_{p}^{2}}\Big{\{}2\mathcal{V}_{m}\mathcal{V}_{e}h_{\text{L}}(3\mathcal{H}^{2}-a^{\prime\prime}/a)-i\mathcal{H}\big{(}(5\mathcal{V}_{m}^{2}+3\mathcal{V}_{e}^{2})\partial_{z}h_{\text{L}}+2i\mathcal{V}_{m}\mathcal{V}_{e}h_{\text{L}}^{\prime}\big{)}+i(\mathcal{V}_{m}^{2}+\mathcal{V}_{e}^{2})\partial_{z}h_{\text{L}}^{\prime}\Big{\}}=0,\end{split}$ (30)
$h_{\text{R}}^{\prime\prime}+2\mathcal{H}h_{\text{R}}^{\prime}-\partial^{2}_{z}h_{\text{R}}-\frac{\xi}{m_{p}^{2}}\Big{\{}2\mathcal{V}_{m}\mathcal{V}_{e}h_{\text{R}}(3\mathcal{H}^{2}-a^{\prime\prime}/a)+i\mathcal{H}\big{(}(5\mathcal{V}_{m}^{2}+3\mathcal{V}_{e}^{2})\partial_{z}h_{\text{R}}-2i\mathcal{V}_{m}\mathcal{V}_{e}h_{\text{R}}^{\prime}\big{)}-i(\mathcal{V}_{m}^{2}+\mathcal{V}_{e}^{2})\partial_{z}h_{\text{R}}^{\prime}\Big{\}}=0.$ (31)
Due to the presence of the Riemann tensor at linear order, note that $a^{\prime\prime}$ appears. In order to cast the (momentum space) equations in the form of (29), we need to remove the dependence on $a^{\prime\prime}$. We can do this by rescaling the amplitudes as $\tilde{h}_{\text{R,L}}=a^{p}h_{\text{R,L}},$ then solve for $p$. It is tedious but straightforward to show that $p=2\xi\mathcal{V}_{m}\mathcal{V}_{e}/m_{p}^{2}$, and now writing $\tilde{h}_{\text{R,L}}(z,\eta)=e^{ikz}\bar{h}_{\text{R,L}}(\eta),$ we obtain
$\bar{h}_{\text{L}}^{\prime\prime}+\Big{\{}2\mathcal{H}-\big{[}A\mathcal{H}+Bk\big{]}\Big{\}}\bar{h}_{\text{L}}^{\prime}+\Big{\{}k^{2}-\big{[}Ck\mathcal{H}+D\mathcal{H}^{2}\big{]}\Big{\}}\bar{h}_{\text{L}}=0,$ (32)
$\bar{h}_{\text{R}}^{\prime\prime}+\Big{\{}2\mathcal{H}+\big{[}\tilde{A}\mathcal{H}+\tilde{B}k\big{]}\Big{\}}\bar{h}_{\text{R}}^{\prime}+\Big{\{}k^{2}+\big{[}\tilde{C}k\mathcal{H}+\tilde{D}\mathcal{H}^{2}\big{]}\Big{\}}\bar{h}_{\text{R}}=0,$ (33)
where the constants are respectively
$\begin{split}A&=-6\xi\tilde{\mathcal{V}}_{m}\tilde{\mathcal{V}}_{e},\\ B&=\xi(\tilde{\mathcal{V}}_{m}^{2}+\tilde{\mathcal{V}}_{e}^{2}),\\ C&=2\xi^{2}\tilde{\mathcal{V}}_{m}\tilde{\mathcal{V}}_{e}(\tilde{\mathcal{V}}_{m}^{2}+\tilde{\mathcal{V}}_{e}^{2})-\xi(5\tilde{\mathcal{V}}_{m}^{2}+3\tilde{\mathcal{V}}_{e}^{2}),\\ D&=-8\xi\tilde{\mathcal{V}}_{m}\tilde{\mathcal{V}}_{e}(1+\tilde{\mathcal{V}}_{m}\tilde{\mathcal{V}}_{e}),\end{split}$ (34)
defining dimensionless quantities $\tilde{\mathcal{V}}_{m}=\mathcal{V}_{m}/m_{p},\ \tilde{\mathcal{V}}_{e}=\mathcal{V}_{e}/m_{p}$, and
$\tilde{A}=-A,\ \ \ \tilde{B}=B,\ \ \ \tilde{C}=C,\ \ \ \tilde{D}=-D.$ (35)
From the parameterization (29), two terms can be straightforwardly identified:
$\begin{split}\beta_{1}=B&=\xi(\tilde{\mathcal{V}}_{m}^{2}+\tilde{\mathcal{V}}_{e}^{2}),\\ \gamma_{0}=C&=2\xi^{2}\tilde{\mathcal{V}}_{m}\tilde{\mathcal{V}}_{e}(\tilde{\mathcal{V}}_{m}^{2}+\tilde{\mathcal{V}}_{e}^{2})-\xi(5\tilde{\mathcal{V}}_{m}^{2}+3\tilde{\mathcal{V}}_{e}^{2}),\end{split}$ (36)
however the other two terms, $A$ and $D$, (a) contribute the same sign to both helicity amplitudes and (b) do not clearly fall into the form of (29). For an arbitrary $\mathcal{V}_{m},\mathcal{V}_{e},$ we see that there are parity preserving and parity violating contributions to the EOM.
As is discussed in Altschul:2009ae , it is consistent to set the “electric” part of $B_{\mu\nu}$ equal to zero, which in our notation, is $\mathcal{V}_{e}=0.$ In this simplified scenario, the $a^{\prime\prime}$ term in in (30), (31) vanishes, as well as $A=\tilde{A}=D=\tilde{D}=0$. The transformation to remove $a^{\prime\prime}$ is then trivial (i.e. $p=0$), and the equations of motion become
$\bar{h}_{\text{R,L}}^{\prime\prime}+\Big{\{}2\mathcal{H}+\lambda_{\text{R,L}}\bar{B}k\Big{\}}\bar{h}_{\text{R,L}}^{\prime}+\Big{\{}k^{2}+\lambda_{\text{R,L}}\bar{C}k\mathcal{H}\Big{\}}\bar{h}_{\text{R,L}}=0,$ (37)
such that we can easily map our model parameters to those of Jenks:2023pmk :
$\bar{B}=\xi\tilde{\mathcal{V}}_{m}^{2}=\beta_{1},\ \ \ \ \bar{C}=-5\xi\tilde{\mathcal{V}}_{m}^{2}=\gamma_{0}.$ (38)
This scenario has the additional interesting feature that the interaction (6) does not affect the background cosmology. This is apparent from the form of the interaction’s contribution to the Einstein equations;
$\Theta_{\mu\nu}=\frac{\mathcal{V}_{e}\mathcal{V}_{m}(a^{\prime\prime}/a-4\mathcal{H}^{2})}{a^{6}}\begin{pmatrix}\frac{-4a^{2}\mathcal{H}^{2}}{a^{\prime\prime}/a-4\mathcal{H}^{2}}&&&\\ &2&&\\ &&1&\\ &&&1\end{pmatrix},$ (39)
which clearly vanishes when $\mathcal{V}_{e}=0.$ The cosmological evolution is thus driven by standard matter and the free energy of the KR field, with stress tensor
$T_{\mu\nu}^{\text{B}}=-2\mathcal{V}_{m}^{2}\mathcal{H}^{2}\begin{pmatrix}1&&&\\ &1&&\\ &&-1&\\ &&&-1\end{pmatrix}-a^{2}V(B)\begin{pmatrix}-1&&&\\ &1&&\\ &&1&\\ &&&1\end{pmatrix},$ (40)
independent of the value of $\mathcal{V}_{e}$.
@@
[…]
The $n_{0},m_{0}$ subscript on $\alpha,\beta,\gamma,\delta$ denote their late time, slowly varying approximations. Our solution has $\beta_{1}=-\tfrac{1}{5}\gamma_{0}=\xi\tilde{\mathcal{V}}_{m}^{2},$ thus only $m=0,$ $n=1$ are nontrivial. In the exponent, we therefore have $z_{0}=\text{ln}(1+z)$ and $D_{2}=(1+z)^{-1}D_{C}\equiv D_{A},$ where $D_{C}$ and $D_{A}$ are respectively the comoving and angular-diameter distances. Our solution is then
$h_{\text{R,L}}=h^{\text{GR}}_{\text{R,L}}\text{exp}\Big{\{}\mp\tfrac{\xi\tilde{\mathcal{V}}_{m}^{2}}{2}\Big{(}k(1+z)D_{2}+5iz_{0}\Big{)}\Big{\}}.$ (43)
As we discussed in Sec. I, the interaction being a dimension-four operator implies that the coupling $\xi$ is dimensionless, such that there is not a mass scale setting a clear UV cutoff. This manifests in (43) by way of only $n=1,m=0$ contributions from (37) along with $\alpha_{n}=\delta_{m}=0$; no factors of $\Lambda_{\text{PV}}$ survive in the parameterization (29).
@@
$\tilde{\lambda}_{\text{R,L}}=-\lambda_{\text{R,L}}(5\xi\tilde{\mathcal{V}}_{m}^{2}\mathcal{H}).$ (45)
@@
$\begin{split}v_{g}&=(1+\tfrac{1}{2}\tilde{\lambda}_{\text{R,L}}k^{-1})(1-\tfrac{1}{2}\tilde{\lambda}_{\text{R,L}}k^{-1}+\tfrac{3}{8}\tilde{\lambda}_{\text{R,L}}^{2}k^{-2}+\cdots)\\ &\simeq 1+\tfrac{3}{8}\tilde{\lambda}_{\text{R,L}}^{2}k^{-2}+O\Big{(}\tilde{\lambda}_{\text{R,L}}^{3}k^{-3}\Big{)}\end{split}$ (47)
for $(\xi\tilde{\mathcal{V}}_{m}^{2}\mathcal{H})\ll 1.$ Since the correction appears at $O(\tilde{\lambda}^{2})$, the sign difference between the left and right handed amplitudes squares away such that they propagate with equal group velocity to leading order. And because the correction is non-negative, the model predicts a propagation speed slightly larger than one. Comparing to the phase velocity, we see that

### v2

@@ Abstract
Gravitational wave physics can probe theories that extend beyond General Relativity. Motivated by the attention on the Kalb-Ramond field as a dark matter candidate, in this work, we study a parity violating dimension four operator which couples the dual Riemann curvature to the 2-form field. After mapping the equations of motion for the right- and left-handed gravitational wave amplitudes to the novel parameterization presented in Jenks:2023pmk , we discuss constraints on the model parameters in light of the coincident electromagnetic/gravitational wave signal of GW170817.
@@ I Introduction
Dynamical Chern-Simons theory (dCS) is likely the most studied parity violating model for gravity. In addition to the two usual massless spin-2 degrees of freedom in GR, dCS includes a pseudoscalar field $\phi$ with a canonical kinetic term coupled to the Pontryagin density $\tilde{R}_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, where $\tilde{R}_{\mu\nu\rho\sigma}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}R^{\alpha\beta}_{\ \ \ \rho\sigma}$ is the dual Riemann tensor, which arises in many contexts Yunes:2009hc (see Alexander:2009tp and references therein for a review). Interesting physics follows if the dCS pseudoscalar acts as the inflaton. This includes a preheating mechanism Alexander:2014bsa , leptogenesis Alexander:2004us , as well as the potential for parity violation to be detected in large-scale structure via galaxy correlations Creque-Sarbinowski:2023wmb . Perturbations to rotating black holes were also studied in dCS theory recently in Wagle:2023fwl , and in Nojiri:2019nar , the propagation of gravitational waves in dCS were compared to standard $f(R)$ modified gravity theories. The dCS coupling was also studied in the context of the string axiverse in Yoshida:2017cjl , where the authors considered a late-time scenario where axions are the dominant dark matter component, and discussed constraints on the coupling.
There are numerous other parity violating theories on the market. One such class is the ghost-free higher derivative models Crisostomi:2017ugk ; Qiao:2021fwi , which are characterized by Lagrangians of the form $\mathcal{L}(g,\partial g,\partial^{2}g)$, and individual terms are contractions of the metric, Riemann, and the Levi-Civita tensors. Higher derivative theories are generically plagued by Ostrogradsky instabilities Woodard:2015zca , therefore such theories can at best be valid up to the mass scale of the ghost degrees of freedom from an effective field theory standpoint. Adjacent to these are the ghost-free scalar tensor models, such as those discussed in Nishizawa:2018srh . Here, derivatives of a scalar field couple to contractions of the Riemann and Levi-Civita tensors. As in the higher order derivative theories, constraints on the Lagrangian coefficients must be imposed in order to avoid Ostrogradsky ghosts. There is also a parity violating addition to the Hor̂ava-Lifshitz approach to quantum gravity Horava:2009uw ; Wang:2017brl , which was studied in the context of gravitational wave physics first in Takahashi:2009wc and later in Zhu:2013fja . A different construction was proposed in Takahashi:2022mew ; here, the authors employed conformal transformations of the metric where the conformal factor is a function of the Pontryagin density. This produces a parity violating effect in the Einstein frame, which was recently used to study modifications of the Kerr metric Tahara:2023pyg .
@@
where the right- and left-handed strains are related to the usual $\times$ amplitudes by $h_{\text{R,L}}=\frac{1}{\sqrt{2}}(h_{+}\pm ih_{\times}).$ Varying the Einstein-Hilbert Lagrangian with respect to each amplitude results in simple wave equations for each. Parity violation will then follow provided an additional interaction term sources the left- and right-handed amplitudes with a different sign, say
@@
Parity violating effects in the gravitational sector generically demand a non-minimal coupling to matter, such as the pseudoscalar in dCS theory. It is interesting to entertain the possibility that such matter belongs to a dark sector. Very recently, the Kalb-Ramond (KR) two-form field $B_{\mu\nu}$ was studied as a dark matter candidate Capanelli:2023uwv . The authors showed that the KR field and Kalb-Ramond-like-particles (KRLPs) can be produced via freeze-in and freeze-out mechanisms and can account for the relic density of dark matter today. Others have considered the KR field as a portal between the SM and dark sectors Dashko:2018dsw ; Dick:2020eop . It has also been studied in the context of gravitational lensing Chakraborty:2016lxo , stellar physics Chakraborty:2017uku , and more generally in early universe scenarios BeltranAlmeida:2018nin ; Almeida:2019xzt including leptogenesis deCesare:2014dga , as well as late time cosmic acceleration Orjuela-Quintana:2022jrg ; BeltranAlmeida:2019fou .
@@ II Features of the KR Lagrangian
Thus although $A_{\mu}$ has four components, only three are independent. The subsidiary gauge transformation (10) is a shift involving an exact 1-form. As briefly mentioned in Sec. I, the gauge invariant theory is dual to the KR axion $\varphi$ via $\partial_{\mu}\varphi\sim\varepsilon_{\mu\nu\rho\sigma}H^{\nu\rho\sigma}$ Svrcek:2006yi ; Smailagic:2001ch and as is usual in axion physics, couplings to matter fields are generically parity violating. The situation is different here, however, due to presence of the potential and the non-minimal coupling to the dual Riemann tensor.
The gauge invariance (9), (10) being broken explicitly by the potential $V$ and the interaction (6) has analogies to the Proca theory of a massive vector boson, which is known to be dual to the massive KR field in certain regimes Hell:2021wzm ; Capanelli:2023uwv . Regarding interactions with the gravitational waves, we will see that a simple and interesting scenario occurs when the KR field’s magnitude settles at a non-zero, constant (classical) value, $B^{2}=B_{(0)}^{2}$. This can be modeled by a potential of the form
@@
where $\bar{g}^{\mu\nu}$ is the unperturbed metric. The explicit form of the potential (11) is not relevant for this work, beyond requiring that it possesses a nontrivial stable minimum [^4] A potential of this form can arise rather naturally. In Assuncao:2019azw , it was shown that an effective potential for the KR field possessing a minimum emerges from quantum corrections following couplings to fermions. . In theories of this type, such potentials have the consequence that the physics can display (spontaneous) Lorentz violation, despite the Lagrange density being Lorentz invariant Altschul:2009ae .
@@ V GW solution in FLRW background
where primes denote partial derivatives with respect to conformal time $\eta$, $\mathcal{H}\equiv a^{\prime}/a$ is the Hubble parameter, $\lambda_{\text{R,L}}=\pm 1,$ and $\Lambda_{\text{PV}}$ is the EFT cutoff. The summation indices are such that $n$ runs over odd integers and $m$ runs over even integers, and the $\{\alpha_{n},\beta_{n},\gamma_{m},\delta_{m}\}$ functions generically depend on the conformal time and are theory dependent [^6] The parameterization written down in Jenks:2023pmk focuses solely on parity odd effects. Recently, in Daniel:2024lev , the parameterization was extended to include parity even effects, with a focus on Chern-Simons and the Gauss-Bonnet coupling. Similar models were also studied in Satoh:2007gn ; Satoh:2008ck in the context of string-inspired inflationary scenarios.. For metric fluctuations on top of an FLRW cosmology in conformal coordinates, the equations of motion for the gravitational wave amplitudes in our model become
$\begin{split}\Big(h_{\text{R,L}}^{\prime\prime}+2\mathcal{H}h_{\text{R,L}}^{\prime}-\partial^{2}_{z}h_{\text{R,L}}\Big)\Big(m_{p}^{2}-3\xi\mathcal{V}_{e}\mathcal{V}_{m}\Big)+i\lambda_{\text{R,L}}v^{2}\partial_{z}h_{\text{R,L}}=0,\end{split}$ (30)
where $v^{2}=\mathcal{V}_{m}^{2}-\mathcal{V}_{e}^{2}$ is the minimum of the potential. Passing to momentum space by writing $\tilde{h}_{\text{R,L}}(z,\eta)=e^{ikz}\bar{h}_{\text{R,L}}(\eta)$ and simplifying, we obtain
$h_{\text{R,L}}^{\prime\prime}+2\mathcal{H}h_{\text{R,L}}^{\prime}+k^{2}\Big(1-\lambda_{\text{R,L}}\frac{\mathcal{H}}{k}\frac{\xi\tilde{v}^{2}}{1-3\xi\tilde{\mathcal{V}}_{e}\tilde{\mathcal{V}}_{m}}\Big)h_{\text{R,L}}=0,$ (31)
where $\tilde{v}=v/m_{p}$, $\tilde{\mathcal{V}}_{e}=\mathcal{V}_{e}/m_{p}$, and $\tilde{\mathcal{V}}_{m}=\mathcal{V}_{m}/m_{p}$. From the parameterization (29), we can immediately identify
$\begin{split}\gamma_{0}=-\frac{\xi\tilde{v}^{2}}{1-3\xi\tilde{\mathcal{V}}_{e}\tilde{\mathcal{V}}_{m}}\approx-\xi\tilde{v}^{2},\end{split}$ (32)
since $\tilde{\mathcal{V}}_{e},\tilde{\mathcal{V}}_{m}\ll 1$. We see that the interaction produces solely birefringence effects in the phase of the gravitational wave polarizations at lowest order.
@@
$\begin{split}D_{\alpha}&=(1+z)^{1-\alpha}\int\frac{(1+z)^{\alpha-2}}{H(z)}dz,\\ z_{\alpha}&=(1+z)^{-\alpha}\int\frac{dz}{(1+z)^{1-\alpha}}.\end{split}$ (34)
The $n_{0},m_{0}$ subscript on $\alpha,\beta,\gamma,\delta$ denote their late time, slowly varying approximations. Our solution has $\gamma_{0}=-\xi\tilde{v}^{2},$ thus only $m=0$ is nontrivial. In the exponent, we therefore have $z_{0}=\text{ln}(1+z)$, and our solution is then
$h_{\text{R,L}}=h^{\text{GR}}_{\text{R,L}}\text{exp}\Big(\mp\frac{i}{2}\xi\tilde{v}^{2}z_{0}\Big).$ (35)
As we discussed in Sec. I, the interaction being a dimension-four operator implies that the coupling $\xi$ is dimensionless, such that there is not a mass scale setting a clear UV cutoff. This manifests in (35) by way of only an $m=0$ contribution from (33) along with $\alpha_{n}=\delta_{m}=0$; no factors of $\Lambda_{\text{PV}}$ survive in the parameterization (29), although a dimension-four operator in principle could produce $\beta_{1}\neq 0.$
@@
$\tilde{\lambda}_{\text{R,L}}=-\lambda_{\text{R,L}}(\xi\tilde{v}^{2}\mathcal{H}).$ (37)
@@
$\begin{split}v_{g}&=(1+\tfrac{1}{2}\tilde{\lambda}_{\text{R,L}}k^{-1})(1-\tfrac{1}{2}\tilde{\lambda}_{\text{R,L}}k^{-1}+\tfrac{3}{8}\tilde{\lambda}_{\text{R,L}}^{2}k^{-2}+\cdots)\\ &\simeq 1+\tfrac{1}{8}\tilde{\lambda}_{\text{R,L}}^{2}k^{-2}+O\Big(\tilde{\lambda}_{\text{R,L}}^{3}k^{-3}\Big)\end{split}$ (39)
for $(\xi\tilde{v}^{2}\mathcal{H})\ll 1.$ Since the correction appears at $O(\tilde{\lambda}^{2})$, the sign difference between the left and right handed amplitudes squares away such that they propagate with equal group velocity to leading order. And because the correction is non-negative, the model predicts a propagation speed slightly larger than one. Comparing to the phase velocity, we see that